Hitting the baseball over the lights
Meanwhile at the ballpark Juan was practicing his hitting while talking to the girls
standing around watching. He started telling them how he hit the ball over the 50 foot
light tower yesterday. The girls were unsure that he really did that.
The formula for this hit is: H(x)=-16x^2+60x+4 where h is the height of the ball and x is the number of seconds the ball is in the air.
How high did Juan actually hit the ball?

Answer by rothauserc(4718)

H(x) = -16x^2 +60x +4
:
This is a parabola that opens downward
:
The x value for the max height is given by the following formula
:
x = -b/2a = -60 / (2*(-16) = 1.875
:
now substitute for x in the H(x) formula
:
-16*(1.875)^2 + 60*(1.875) + 4 = 60.25
:
Juan hit the ball 60.25 feet high
: